# Active Inference in Contextual Multi-Armed Bandits for Autonomous Robotic Exploration

### v1

@@ I Introduction
For instance, consider a scenario, as illustrated in Fig. 1, where an aerial robot must autonomously select the most promising site for sampling a desired mineral rock specimen during a follow-up in-situ survey mission [5]. This search site selection is based on remote sensing data (e.g. hyperspectral information) gathered from predetermined search sites using a lightweight sensor (e.g. spectrometer). In this scenario, however, the sensing returns obtained by directing sensors are stochastic for various reasons, for example, due to observation noise and the variability in the targeted coordinates for each site sampling instance. Moreover, model parameters used to predict the likelihood of observing each outcome (e.g. a detected mineral specimen) based on these returns are not known beforehand. Thus, the robot needs to carefully strike a balance between exploitation (increasing the certainty in sites where desired mineral specimens are expected) and exploration (decreasing the uncertainty in poorly explored search sites). This kind of sequential decision-making problem can be formulated as a contextual multi-armed bandit (CMAB), which has been widely studied in recommendation systems, finance, and healthcare [6, 7, 8] [^1] This sequential decision making problem can also be formulated using other mathematical frameworks such as POMDPs [9, 10]. However, these frameworks require careful definitions of state transition and reward functions. Especially, in new and uncertain environments, reward tuning is tricky and time-consuming. Also, these do not incorporate human expert preference regarding outcomes easily. Thus, instead, we opt to study simpler CMAB formulation, which allows us to advantageously abstract certain lower-level behavioral aspects of the data site selection problem..
@@
However, in general, bandit problems–depending on their scale and complexity–often require a large number of iterative interactions with the environment to finalize the optimal option. This need for numerous iterations can become a bottleneck when applying this mathematical framework to practical robotic applications, such as space exploration and mineralogical surveys on Earth, where resources and time are often constrained. Moreover, existing conventional methods in CMABs typically do not explicitly take into account human experts’ (e.g. scientists’) prior preferences regarding outcomes in their decision-making processes. As a consequence, the decisions derived from these methods may not always align with what humans are actually interested in observing, leading to the decrease in mission efficiency. In light of these backdrops, our previous studies [11, 12] sought to emulate the approach taken by astronaut Harrison Schmitt during the Apollo 17 mission, where he combined in-situ findings with geological expert knowledge to advance lunar geology [13]. To achieve a similar behavior in robotic systems, we applied active inference (AIF) [14, 15, 16]–which originated in computational neuroscience and has recently gained traction in robotics–to develop option selection strategies for CMABs that are informed by expert-provided prior preferences on observations. While these studies showed that AIF agents could efficiently identify the best option for humans compared to other strategies when expert prior preference is stationary, the contextual information used for decision-making and the true hidden model parameters associated with the options were randomly generated, which does not reflect real-world conditions. Hence, in this article, we aim to validate the applicability of the AIF-based option selection methodology in realistic problem scenarios. Specifically, the key contributions and novelty with respect to previous studies [11, 12] are:
@@
  • Demonstrating the effectiveness of the AIF-based option selection algorithm using real scientific data, namely based on hyperspectral data collected by AVIRIS-NG [17] and mineral label data created by geologists [18].
@@ II-B Free Energy Principle and Active Inference
Active inference (AIF) is a mathematical framework that applies the FEP specifically to the behavioral norms of biological agents [14, 15]. In the field of neuroscience, it has been used to understand the characteristic behavioral mechanisms observed in patients with autism [29, 30]. Recently, it has also garnered attention in the field of robotics for state estimation, adaptive control, and for decision making under uncertainty [31, 32, 16, 33]. The reason for this lies in the expected free energy (EFE) objective function characterizing AIF. Although a detailed explanation is provided in Sec. III, the EFE for each possible option/action in the MAB/CMAB context consists of the (negative) value resulting from selecting a particular option and the (negative) information gain representing how much the uncertainty about hidden states is reduced by taking that option. Consequently, by optimizing (i.e. minimizing) the EFE, agents can naturally take an action balancing exploitation and exploration. Additionally, since preference information regarding outcome observations, known as prior preference, can be externally incorporated into the value term, agents take actions biased towards obtaining desired observations. This characteristic has recently been studied for Pareto point selection problems in multi-objective reinforcement learning[34].
@@ III-A Problem Statement
Suppose the total number of options (e.g. search sites) taken into account by a robot is $K\in\mathbb{N}$. Note that these options are equivalent to the bandit arms and selecting an option $k\in\{1,\cdots,K\}$ is denoted as $a=k$ (for the ease of notation, in the following, we use $a_{k}\leftrightarrow a=k$). Additionally, suppose that a semantic observation $o_{k}$ (e.g. mineral label) of each option $k$ from an observation source is multicategorical across $F$ labels, i.e. $o_{k}=f,f\in\mathcal{F}=\{1,\cdots,F\}$. Therefore, the probability that a feature $f$ is observed by investigating an option $k$ at a decision instance $t$ can be described as the following softmax likelihood function [35, 36],
@@ III-B2 Special Case: Multivariate Gaussian Prior and Softmax Observation Likelihood
[…]
@@
where H is the Hessian. Thus, if we define $A=-$H and by taking the logarithm from (11),
@@
[…]
@@ IV Simulation Study
To verify whether the proposed active inference option selection method is effective not only for CMABs formulated with randomly generated hidden parameters and contexts, as in previous studies [11, 12], but also for CMABs formulated based on actual scientific data, a mineral search site selection study is considered. In the following subsections, we begin with an overview of the motivating autonomous robotic exploration scenario focusing on surface mineralogical surveys. We then describe the hyperspectral and mineral label dataset used as contexts and outcome observations. This is followed by an explanation of the preprocessing steps and the result of learning the true hidden parameters necessary for calculating the cumulative regret. Finally, we detail the simulation setup and present the results of Monte Carlo simulation experiments under both static and dynamic human prior preferences.
@@ IV-C Training True Latent Parameters
After these preprocessing steps, as shown in Fig. 5, $K$ non-overlapping search sites are selected, each with dimensions of $200$ pixels in width and $250$ pixels in height, and the pairs of hyperspectral and mineral label data are combined as datasets. To train the true best-fit softmax parameters $\vec{\Theta}^{*}_{k}$, the dataset is split into training (80%) and test sets (20%) and the softmax regression (i.e. multinomial logistic regression) with the Adam optimizer [47] is performed with PyTorch [48]. The average accuracy (i.e. the proportion of correctly classified samples out of the total number of samples) over all search sites is $76.7\%$ (note that min/max accuracy is $62.3\%$ and $93.1\%$, respectively). Despite having a relatively low accuracy as a classifier, it effectively captures the noise present in the measurement process as shown in Fig. 6. This is further confirmed by examining the histograms of the learned bias values. As depicted in Fig. 7, each subplot exhibits significant negative values (around $-20$). This observation indicates that the trained classifier discerns the absence of certain minerals in these search sites. Additionally, considering that other research utilizing a similar dataset also demonstrates comparable accuracy values, it suggests that this classifier is satisfactory to provide a best fit baseline comparison [43].

### v2

@@ I Introduction
For instance, consider a scenario, as illustrated in Fig. 1, where an aerial robot must autonomously select the most promising site for sampling a desired mineral rock specimen during a follow-up in-situ survey mission [5]. This search site selection is based on remote sensing data (e.g. hyperspectral information) gathered from predetermined search sites using a lightweight sensor (e.g. spectrometer). In this scenario, however, the sensing returns obtained by directing sensors are stochastic for various reasons, for example, due to observation noise and the variability in the targeted coordinates for each site sampling instance. Moreover, model parameters used to predict the likelihood of observing each outcome (e.g. a detected mineral specimen) based on these returns are not known beforehand. Thus, the robot needs to carefully strike a balance between exploitation (increasing the certainty in sites where desired mineral specimens are expected) and exploration (decreasing the uncertainty in poorly explored search sites). This sequential decision-making problem can be formulated with mathematical frameworks such as partially observable Markov decision processes (POMDPs) [6, 7] and Gaussian processes (GPs) [8, 9]. In the case of POMDPs, however, careful definitions of a state transition function, a reward function, and planner hyperparameters (e.g. a planning depth and a discount factor) are required. Reward and hyperparameter tuning, in particular, can be tricky and time-consuming in new and uncertain environments [10]. On the other hand, in the case of GPs, while they can leverage a spatial structure of an environment, the computation cost of kernel functions is significant for real-time operation [11, 12]. Additionally, neither framework easily incorporates human expert preferences regarding outcomes easily. Thus, instead, we opt to study simpler contextual multi-armed bandit (CMAB) formulation, which has been widely studied in recommendation systems, finance, healthcare, and recently, robotics [13, 14, 15, 16], and allows us to advantageously abstract certain lower-level behavioral aspects of the search site selection problem.
@@
However, in general, bandit problems–depending on their scale and complexity–often require a large number of iterative interactions with the environment to finalize the optimal option. This need for numerous iterations can become a bottleneck when applying this mathematical framework to practical robotic applications, such as space exploration and mineralogical surveys on Earth, where resources and time are often constrained. Moreover, existing conventional methods in CMABs typically do not explicitly take into account human experts’ (e.g. scientists’) prior preferences regarding outcomes in their decision-making processes. As a consequence, the decisions derived from these methods may not always align with what humans are actually interested in observing, leading to the decrease in mission efficiency. In light of these backdrops, our previous studies [17, 18] sought to emulate the approach taken by astronaut Harrison Schmitt during the Apollo 17 mission, where he combined in-situ findings with geological expert knowledge to advance lunar geology [19]. To achieve a similar behavior in robotic systems, we applied active inference (AIF) [20, 21, 22]–which originated in computational neuroscience and has recently gained traction in robotics–to develop option selection strategies for stationary, independent, and linear CMABs that are informed by expert-provided prior preferences on observations. While these studies showed that AIF agents could efficiently identify the best option for humans compared to other strategies when expert prior preference is stationary, the contextual information used for decision-making and the true hidden model parameters associated with the options were randomly generated, which does not reflect real-world conditions. Hence, in this article, we aim to validate the applicability of the AIF-based option selection methodology in realistic problem scenarios. Specifically, the key contributions and novelty with respect to previous studies [17, 18] are:
@@
  • Demonstrating the effectiveness of the AIF-based option selection algorithm using real scientific data, namely based on hyperspectral data collected by AVIRIS-NG [23] and mineral label data created by geologists [24]. This study marks the first application of active inference in geological data exploration.
@@ II-B Free Energy Principle and Active Inference
Active inference (AIF) is a mathematical framework that applies the FEP specifically to the behavioral norms of biological agents [20, 21]. In the field of neuroscience, it has been used to understand the characteristic behavioral mechanisms observed in patients with autism [35, 36]. Recently, it has also garnered attention in the field of robotics for state estimation, adaptive control, and for decision making under uncertainty [37, 38, 22, 39]. The reason for this lies in the expected free energy (EFE) objective function characterizing AIF. Although a detailed explanation is provided in Sec. III, the EFE for each possible option/action in the MAB/CMAB context consists of the (negative) value resulting from selecting a particular option and the (negative) information gain (i.e. mutual information commonly known in robotics [40, 41] and Bayesian experimental design [42, 43]) representing how much the uncertainty about hidden states is reduced by taking that option. Consequently, by optimizing (i.e. minimizing) the EFE, agents can naturally take an action balancing exploitation and exploration. Additionally, since preference information regarding outcome observations, known as prior preference, can be externally incorporated into the value term, agents take actions biased towards obtaining desired observations. This characteristic has recently been studied for Pareto point selection problems in multi-objective reinforcement learning[44].
@@ III-A Problem Statement
Suppose the total number of options (e.g. search sites) taken into account by a robot is $K\in\mathbb{N}$. Note that these options are equivalent to the bandit arms and selecting an option $k\in\{1,\cdots,K\}$ is denoted as $a=k$ (for the ease of notation, in the following, we use $a_{k}\leftrightarrow a=k$). Additionally, suppose that a semantic observation $o_{k}$ (e.g. mineral label) of each option $k$ from an observation source is multicategorical across $F$ labels, i.e. $o_{k}=f,f\in\mathcal{F}=\{1,\cdots,F\}$. Therefore, the probability that a feature $f$ is observed by investigating an option $k$ at a decision instance $t$ can be described as the following softmax likelihood function [45, 46] [^1] It is also natural to choose a Dirichlet distribution as a prior and a categorical distribution as an observation likelihood, as their conjugacy allows for easy posterior calculation [45]. Nevertheless, this approach cannot easily incorporate continuous contextual information (such as hyperspectral data) associated with the options. Hence, the softmax function, which is one of hybrid discrete-continuous likelihood functions and has gained attention in the field of multi-sensor fusion [47, 48, 49, 50], is adopted.,
@@ III-B2 Special Case: Multivariate Gaussian Prior and Softmax Observation Likelihood
$\frac{1}{2}\Big{\{}\sum_{r=1}^{(C+1)\times F}$ $(\Theta_{k,r}-\Theta_{k,r}^{(0)})\frac{\partial\log g(\vec{\Theta}_{k}^{(0)})}{\partial\Theta_{k,r}}\Big{\}}^{2}$
$=\frac{1}{2}\Big{\{}(\vec{\Theta}_{k}-\vec{\Theta}_{k}^{(0)})^{T}\nabla\log g(\vec{\Theta}_{k}^{(0)})\Big{\}}^{2},$ (10)
@@
where H is the Hessian (Note that the Hessian can be calculated by computing the Jacobian of the gradient, and the gradient can be derived with an optimizer implemented in the standard scientific computing library such as scipy.optimizer). Thus, if we define $A=-$H and by taking the logarithm from (11),
@@
$q(o|a_{k})=g(\vec{\Theta}_{k,MAP})\cdot\frac{(2\pi)^{\frac{(C+1)\times F}{2}}}{|A|^{\frac{1}{2}}}.$ (13)
@@ IV Simulation Study
To verify whether the proposed active inference option selection method is effective not only for stationary, independent, and linear CMABs formulated with randomly generated hidden parameters and contexts, as in previous studies [17, 18], but also for CMABs formulated based on actual scientific data, a mineral search site selection study is considered. In the following subsections, we begin with an overview of the motivating autonomous robotic exploration scenario focusing on surface mineralogical surveys. We then describe the hyperspectral and mineral label dataset used as contexts and outcome observations. This is followed by an explanation of the preprocessing steps and the result of learning the true hidden parameters necessary for calculating the cumulative regret. Finally, we detail the simulation setup and present the results of Monte Carlo simulation experiments under both static and dynamic human prior preferences.
@@ IV-C Training True Latent Parameters
After these preprocessing steps, as shown in Fig. 5, $K$ non-overlapping search sites are selected, each with dimensions of $200$ pixels in width and $250$ pixels in height, and the pairs of hyperspectral and mineral label data are combined as datasets. To train the true best-fit softmax parameters $\vec{\Theta}^{*}_{k}$, the dataset is split into training (80%) and test sets (20%) and the softmax regression (i.e. multinomial logistic regression) with the Adam optimizer [59] is performed with PyTorch [60]. The average accuracy (i.e. the proportion of correctly classified samples out of the total number of samples) over all search sites is $76.7\%$ (note that min/max accuracy is $62.3\%$ and $93.1\%$, respectively). Despite having a relatively low accuracy as a classifier, it effectively captures the noise present in the measurement process as shown in Fig. 6. This is further confirmed by examining the histograms of the learned bias values. As depicted in Fig. 7, each subplot exhibits significant negative values (around $-20$). This observation indicates that the trained classifier discerns the absence of certain minerals in these search sites [^4] For example, in Fig. 6, it can be observed that minerals of Classes $2$ and $13$ are absent at sites $3$ and $6$. As the number of such absent minerals increases, the frequency of large negative values in Fig. 7 also increases.. Additionally, considering that other research utilizing a similar dataset also demonstrates comparable accuracy values, it suggests that this classifier is satisfactory to provide a best fit baseline comparison [56].